13a+5b=-11
13a+11b=7
how do you do this problem in elimination?

Answer by TimothyLamb(4379)

---
13a+5b=-11
13a+11b=7
---
put the system of linear equations into standard form
---
13x + 5y = -11
13x + 11y = 7
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = a = -2
y = b = 3
---
system matrix
13 5
13 11
---
inverse of system matrix
0.14102564 -0.064102564
-0.16666667 0.16666667
---
determinant of system matrix = 78
---
now solution by elimination method:
---
(13a + 5b = -11)-1
13a + 11b = 7
---
-13a - 5b = 11
13a + 11b = 7
---
add the system:
---
6b = 18
---
b = 3
---
13a + 5b = -11
13a + 5(3) = -11
13a = -11 - 15
13a = -26
---
a = -2
---
solution by elimination method:
x = a = -2
y = b = 3
---
NOTE:
solution by matrix method agrees with solution by elimination method.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php